# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ III-C Reconfigurable Scattering Device: Mode Transfer Matrix
where ${\bf{I}}_{K}$ is the identity matrix of size $K$, and ${\bf{R}}={\bf{R}}(\bm{\theta})={\bf{G}}\,{\bf{D}}({\bf{I}}_{K}-{\bf{G}}\,{\bf{D}})^{-1}$ serves as the reflection matrix (or scattering matrix in the electromagnetic community). Specifically, the $(k,i)$-th element of ${\bf{R}}$ represents the reflection coefficient corresponding to the $k$-th component of the impinging electric field and the $i$-th component of the scattered field, as described by the basis set $\{{\mathbf{\Phi}}_{k}({\bf{r}})\}$. Notably, depending on the available technology, the matrix ${\bf{D}}$ might be constrained to take a particular structure. For example, if we exclude the presence of non-reciprocal materials in the scatterer, the response of the scattering device must verify the reciprocity theorem, hence resulting in a symmetric matrix [40]. Another example is the case where the boundary conditions imposed by the reconfigurable scatterer are local, i.e., $\mathsf{D}({\bf{r}},{\bf{s}})=\mathsf{D}({\bf{r}})\,\delta({\bf{r}}-{\bf{s}})$ and hence ${\mathsf{J}}_{\mathrm{s}}({\bf{r}})=\mathsf{D}({\bf{r}})\,{\mathsf{E}}({\bf{r}})$ in (14) [15]. For instance, if we consider the reconfigurable scattering device to be composed of a set of $K$ uncoupled discrete elements each described by the $k$-th function ${\mathbf{\Phi}}_{k}({\bf{r}})$ of the basis set, the $\left\{{\mathbf{\Phi}}_{k}({\bf{r}})\right\}$’s are disjoint functions, and hence matrix ${\bf{D}}$ is restricted to be diagonal as well as the reflection matrix ${\bf{R}}$ being the elements uncoupled by assumption (see Sec. V for examples).
@@
Suppose the communication modes associated with the first link (from the source space to the scattering device) and the second link (from the scattering device to the receiving space) have been derived through the coupled eigenfunction problems (11) by truncating the series to the first $K$ terms, where $K$ is sufficiently large. Denote with ($\left(\left\{{\bm{\varphi}}^{(\text{T})}_{n}({\bf{r}})\right\},\left\{\mathbf{\Psi}^{(\text{T})}_{n}({\bf{r}})\right\},\sigma_{n}^{(\text{T})}\right)$ the $n$-th basis couple and singular value of the first link, for $n=1,2,\ldots,K$, and with ($\left(\left\{{\bm{\varphi}}^{(\text{R})}_{m}({\bf{r}})\right\},\left\{\mathbf{\Psi}^{(\text{R})}_{m}({\bf{r}})\right\},\sigma_{m}^{(\text{R})}\right)$ the $m$-th basis couple and singular value of the second link, for $m=1,2,\ldots,K$. Denote with ${\mathbf{G}_{\text{T}}={\rm diag}\left(\sigma_{1}^{(\text{T})},\sigma_{2}^{(\text{T})},\ldots,\sigma_{K}^{(\text{T})}\right)\in\mathbb{C}^{K\times K}}$ and $\mathbf{G}_{\text{R}}={\rm diag}\left(\sigma_{1}^{(\text{R})},\sigma_{2}^{(\text{R})},\ldots,\sigma_{K}^{(\text{R})}\right)\in\mathbb{C}^{K\times K}$ the coupling matrices between, respectively, the source and device spaces, and the device and receiving spaces. The effect of the scattering device can be fully described by the device operator ${\bf{C}}={\bf{C}}(\bm{\theta})\in\mathbb{C}^{K\times K}$, whose generic element $[{\bf{C}}]_{m,n}$ represents the coupling coefficient between the input communication mode $n$ and the output communication mode $m$. As a consequence, the end-to-end coupling between the source space and the receiving space is given by
${\bf{H}}(\bm{\theta})=\mathbf{G}_{\text{R}}\,{\bf{C}}(\bm{\theta})\,\mathbf{G}_{\text{T}}\,.$ (17)
@@
The device operator ${\bf{C}}(\bm{\theta})$, also known as the mode transfer matrix, is determined by the specific structure and technology of the scattering device, as well as its configuration parameters $\bm{\theta}$, which enable the realization of different end-to-end processing functionalities ${\bf{H}}(\bm{\theta})$. In essence, any linear scattering device can be seen as a mode converter (see Fig. 5) [39].
Figure 5: The reconfigurable scattering device as a mode converter. The spheres represent the input and output basis functions.
@@
We now derive the relationship between ${\bf{C}}(\bm{\theta})$ and ${\bf{D}}(\bm{\theta})$. By defining ${\bf{U}}\in\mathbb{C}^{K\times K}$, with elements $[{\bf{U}}]_{m,k}=\left<{{\bm{\varphi}}^{(\text{R})}_{m}({\bf{r}})}\,,{{\mathbf{\Phi}}_{k}({\bf{r}})}\right>$ as the transformation matrix between the basis $\left\{{\bm{\varphi}}^{(\text{R})}_{m}({\bf{r}})\right\}$ and $\left\{{\mathbf{\Phi}}_{k}({\bf{r}})\right\}$, and ${\bf{V}}\in\mathbb{C}^{K\times K}$, with elements $[{\bf{V}}]_{n,k}=\left<{\mathbf{\Psi}^{(\text{T})}_{n}({\bf{r}})}\,,{{\mathbf{\Phi}}_{k}({\bf{r}})}\right>$, the transformation matrix between the basis $\left\{{\mathbf{\Phi}}_{k}({\bf{r}})\right\}$ and $\left\{\mathbf{\Psi}^{(\text{T})}_{n}({\bf{r}})\right\}$, we can write
${\bf{C}}(\bm{\theta})={\bf{U}}\,{\bf{D}}(\bm{\theta})\,({\bf{I}}_{K}-{\bf{G}}\,{\bf{D}}(\bm{\theta}))^{-1}\,{\bf{V}}^{-1}\,.$ (18)
@@
It is useful to examine how (18) simplifies under the assumption of weak reflections, i.e., when $|{\mathsf{E}}_{\mathrm{s}}({\bf{r}})|\ll|{\mathsf{E}}_{\mathrm{i}}({\bf{r}})|$ for ${\bf{r}}\in{\cal{S}}$, which corresponds to $\|{\bf{G}}\,{\bf{D}}(\bm{\theta})\|\ll 1$. Specifically, we have
${\bf{C}}(\bm{\theta})\approx{\bf{U}}\,{\bf{D}}(\bm{\theta})\,{\bf{V}}^{-1}\,.$ (19)
@@
This corresponds to the Born approximation, which involves using the incident field as the driving field at each point in the scattering device, rather than the total EM field. This approximation is valid when the scattered field is small relative to the incident field on the scattering device. Notably, this result is equivalent to removing the feedback loop in Fig. 4. Depending on the specific technology being analyzed, the use of the approximation (19) instead of (18) should be considered with caution, as it may fail to capture important physical phenomena. An example of this is the Floquet modes in periodic metamaterials, which correspond to spurious reflections in undesired directions [22, 15].
@@
In conclusion, the ESP problem can be framed as finding the optimal configuration parameter set $\bm{\theta}$ that best approximates a given objective end-to-end processing functionality ${\bf{H}_{\text{o}}}$ according to a specific criterion. This can be formulated as the following minimization problem:
$\hat{\bm{\theta}}=\arg\min_{\bm{\theta}}\|\mathbf{G}_{\text{R}}\,{\bf{C}}(\bm{\theta})\,\mathbf{G}_{\text{T}}-{\bf{H}_{\text{o}}}\|_{\mathrm{F}}$ (20)
@@
subject to a constraint on the radiated power, where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm. Depending on the adopted technology, additional constraints should be incorporated into the minimization problem to account for the specific structure and symmetries of the matrix ${\bf{D}}(\bm{\theta})$ and hence ${\bf{C}}(\bm{\theta})$. This underscores the need for tailored optimization algorithms based on physically consistent models. In this regard, specific examples will be provided in Sec. V-B.
@@ IV Fundamental Limits
The ultimate limit in terms of the number of DoF, i.e., the flexibility of the reconfigurable scattering device to produce the desired EM response, can be obtained supposing that through a proper configuration of the device and illumination from the source space, it is possible to generate any induced current distribution on ${\cal{S}}$. In the following, we see that even in the case of an unbounded space (infinite receiver’s size), the number of DoF is limited by the size of the scattering device compared to the wavelength. This number of DoF is further limited when finite source and receiver spaces are considered. In any case, these theoretical limits have to be intended as upper bounds because the particular technology considered might pose further constraints.
@@ IV-A DoF in the Unbounded Space
In the following, we revisit the main results regarding the number of DoF for the radiating EM field in an unbounded space, given a current density ${\mathsf{J}}({\bf{r}})$ defined within a limited space ${\cal{S}}$, using signal processing arguments.
@@ IV-A1 DoF for 1D Structures
In this case, ${\cal{S}}$ is a line segment of length $L$ with coordinate $r_{x}\in[-L/2,L/2]$. Without loss of generality, we consider vertical EM waves polarization. Suppose ${\mathsf{J}}({\bf{r}})=J(r_{x})\,\delta(r_{y})\,\delta(r_{z})\,{\hat{\bf{y}}}$ defined in ${\cal{S}}$. It is ${\widetilde{\mathsf{J}}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{J}}({\bf{r}})\right]=\mathcal{F}\left[J(r_{x})\right]{\hat{\bf{y}}}={\widetilde{J}}(\kappa_{x})\,{\hat{\bf{y}}}={\widetilde{\mathsf{J}}}(\kappa_{x})$, where ${\widetilde{J}}(\kappa_{x})=\mathcal{F}\left[J(r_{x})\right]$. It is well-known that every function $J(r_{x})$ limited in $[-L/2,L/2]$ can be represented as an infinite series expansion of Fourier basis functions (i.e., modes)
$J(r_{x})=\sum_{n=-\infty}^{\infty}j_{n}\cdot\frac{1}{\sqrt{L}}\text{Rect}\left({\frac{r_{x}}{L}}\right)\,e^{\jmath k_{n}\,r_{x}}$ (21)
@@
${\widetilde{J}}(\kappa_{x})=\sum_{n=-\infty}^{\infty}j_{n}\,\sqrt{L}\,\text{Sinc}\left({\frac{L}{2\pi}\left(\kappa_{x}-k_{n}\right)}\right)$ (22)
being $\text{Sinc}\left({x}\right)=\sin(\pi x)/(\pi x)$. Note that, according to (6), only the modes for which $|k_{n}|\leq k_{0}=2\pi/\lambda$ radiate. Then, only $N=2L/\lambda$ terms in (22) fall in the visible region. [^4] More precisely, we assume that $L\gg\lambda$; otherwise, the tails of certain Sinc functions centered in the non-visible region could significantly contribute to the visible region. Being the wavenumber support limited (i.e., ${\widetilde{J}}(\kappa_{x})$ band limited in $[-k_{0},k_{0}]$), we can say that any EM radiated field that can be generated by a current line source of length $L$ can be fully represented by sampling it in the wavenumber domain at frequencies $k_{n}=2n\pi/L$ for $|n|\leq N/2$. Thus, the required number $N$ of samples leads to the DoF in the unbounded space, that is, $DoF=\frac{2L}{\lambda}$ (per polarization). To draw a parallel, when beamforming is operated using $N$ half-wavelength spaced antennas arranged as a uniform linear array (ULA), $N$ orthogonal beams (i.e., $2L/\lambda$ when the antenna spacing is $\lambda/2$) can be realized. These can be adopted to reach users in every possible angular direction (e.g., for initial access). These beams correspond to the set of orthogonal transmission directions that can be spanned by the ULA considering the classical discrete Fourier transform (DFT)-based codebook of beamforming vectors given by
$\frac{1}{\sqrt{N}}\begin{bmatrix}1,\,\,e^{-\frac{\jmath 2\pi n}{N}},\,\,e^{-\frac{\jmath 2\pi 2n}{N}},\,\,\ldots,\,\,e^{-\frac{\jmath 2\pi(N-1)n}{N}}\end{bmatrix}^{{\mathrm{T}}}=\frac{1}{\sqrt{N}}\begin{bmatrix}1,\,\,e^{-\frac{\jmath k_{n}\lambda}{2}},\,\,e^{-\frac{\jmath 2k_{n}\lambda}{2}},\,\,\ldots,\,\,e^{-\frac{\jmath(N-1)k_{n}\lambda}{2}}\end{bmatrix}^{{\mathrm{T}}}\,.$ (23)
The different vectors correspond to the transmission directions towards angles $\gamma_{n}=\arcsin{\frac{2n}{N}}$, for $n=\pm 1,\pm 2,\ldots,\left\lfloor\frac{N}{2}\right\rfloor$. The characteristic of these beams is that each one points towards the null directions of all the other beams. Every possible channel vector (i.e., direction) can be realized as a linear combination of the vectors in (23), which is an orthonormal basis for $\mathbb{C}^{N}$.

### v3

@@ III-C Reconfigurable Scattering Device: Mode Transfer Matrix
where ${\mathbf{I}}_{K}$ is the identity matrix of size $K$, and ${\mathbf{R}}(\bm{\theta})={\mathbf{G}}\,{\mathbf{D}}(\bm{\theta})({\mathbf{I}}_{K}-{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta}))^{-1}$ serves as the reflection matrix (or scattering matrix in the electromagnetic community). Specifically, the $(k,i)$-th element of ${\mathbf{R}}(\bm{\theta})$ represents the reflection coefficient corresponding to the $i$-th component of the impinging electric field and the $k$-th component of the scattered field, as described by the basis set $\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\}$. Notably, depending on the available technology, the matrix ${\mathbf{D}}(\bm{\theta})$ might be constrained to take a particular structure. For example, if we exclude the presence of non-reciprocal materials in the scatterer, the response of the scattering device must verify the reciprocity theorem, hence resulting in a symmetric matrix [39]. Another example is the case where the boundary conditions imposed by the reconfigurable scatterer are local, i.e., $\mathsf{D}({\mathbf{r}},{\mathbf{s}})=\mathsf{D}({\mathbf{r}})\,\delta({\mathbf{r}}-{\mathbf{s}})$ and hence ${\mathsf{J}}_{\mathrm{s}}({\mathbf{r}})=\mathsf{D}({\mathbf{r}})\,{\mathsf{E}}({\mathbf{r}})$ in (15) [15]. For instance, if we consider the reconfigurable scattering device to be composed of a set of $K$ uncoupled discrete elements each described by the $k$-th function ${\mathbf{\Phi}}_{k}({\mathbf{r}})$ of the basis set, the $\left\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\right\}$’s are disjoint functions, and hence matrix ${\mathbf{D}}(\bm{\theta})$ is restricted to be diagonal as well as the reflection matrix ${\mathbf{R}}(\bm{\theta})$ being the elements uncoupled by assumption (see Section V for examples).
@@
Suppose the communication modes associated with the first link (from the source space to the scattering device) and the second link (from the scattering device to the receiving space) have been derived through the coupled eigenfunction problems (12) by truncating the series to the first $K$ terms, where $K$ is sufficiently large. Denote with $\left(\left\{{\bm{\varphi}}^{(\text{T})}_{n}({\mathbf{r}})\right\},\left\{\mathbf{\Psi}^{(\text{T})}_{n}({\mathbf{r}})\right\},\sigma_{n}^{(\text{T})}\right)$ the $n$-th basis couple and singular value of the first link, for $n=1,2,\ldots,K$, and with $\left(\left\{{\bm{\varphi}}^{(\text{R})}_{m}({\mathbf{r}})\right\},\left\{\mathbf{\Psi}^{(\text{R})}_{m}({\mathbf{r}})\right\},\sigma_{m}^{(\text{R})}\right)$ the $m$-th basis couple and singular value of the second link, for $m=1,2,\ldots,K$. Denote with $\mathbf{G}_{\text{T}}={\rm diag}\left(\sigma_{1}^{(\text{T})},\sigma_{2}^{(\text{T})},\ldots,\sigma_{K}^{(\text{T})}\right)\in\mathbb{C}^{K\times K}$} and $\mathbf{G}_{\text{R}}={\rm diag}\left(\sigma_{1}^{(\text{R})},\sigma_{2}^{(\text{R})},\ldots,\sigma_{K}^{(\text{R})}\right)\in\mathbb{C}^{K\times K}$ the coupling matrices between, respectively, the source and device spaces, and the device and receiving spaces. The effect of the scattering device can be fully described by the device operator ${\mathbf{C}}(\bm{\theta})\in\mathbb{C}^{K\times K}$, whose generic element ${\mathbf{C}}(\bm{\theta})]_{m,n}$ represents the coupling coefficient between the input communication mode $n$ and the output communication mode $m$. As a consequence, the end-to-end coupling between the source space and the receiving space is given by
${\mathbf{H}}(\bm{\theta})=\mathbf{G}_{\text{R}}\,{\mathbf{C}}(\bm{\theta})\,\mathbf{G}_{\text{T}}\,.$ (18)
@@
The device operator ${\mathbf{C}}(\bm{\theta})$, also known as the mode transfer matrix, is determined by the specific structure and technology of the scattering device, as well as its configuration parameters $\bm{\theta}$, which enable the realization of different end-to-end processing functionalities ${\mathbf{H}}(\bm{\theta})$. In essence, any linear scattering device can be seen as a mode converter [38].
@@
We now derive the relationship between ${\mathbf{C}}(\bm{\theta})$ and ${\mathbf{D}}(\bm{\theta})$. By defining ${\mathbf{U}}\in\mathbb{C}^{K\times K}$, with elements ${\mathbf{U}}]_{m,k}=\left<{{\bm{\varphi}}^{(\text{R})}_{m}({\mathbf{r}})}\,,{{\mathbf{\Phi}}_{k}({\mathbf{r}})}\right$> as the transformation matrix between the basis $\left\{{\bm{\varphi}}^{(\text{R})}_{m}({\mathbf{r}})\right\}$ and $\left\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\right\}$, and ${\mathbf{V}}\in\mathbb{C}^{K\times K}$, with elements ${\mathbf{V}}]_{n,k}=\left<{\mathbf{\Psi}^{(\text{T})}_{n}({\mathbf{r}})}\,,{{\mathbf{\Phi}}_{k}({\mathbf{r}})}\right$>, the transformation matrix between the basis $\left\{{\mathbf{\Phi}}_{k}({\mathbf{r}})\right\}$ and $\left\{\mathbf{\Psi}^{(\text{T})}_{n}({\mathbf{r}})\right\}$, we can write
${\mathbf{C}}(\bm{\theta})={\mathbf{U}}\,{\mathbf{D}}(\bm{\theta})\,({\mathbf{I}}_{K}-{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta}))^{-1}\,{\mathbf{V}}^{-1}\,.$ (19)
@@
It is useful to examine how (19) simplifies under the assumption of weak reflections, i.e., when $|{\mathsf{E}}_{\mathrm{s}}({\mathbf{r}})|\ll|{\mathsf{E}}_{\mathrm{i}}({\mathbf{r}})|$ for ${\mathbf{r}}\in{\cal{S}}$, which corresponds to $\|{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta})\|\ll 1$. Specifically, we have
${\mathbf{C}}(\bm{\theta})\approx{\mathbf{U}}\,{\mathbf{D}}(\bm{\theta})\,{\mathbf{V}}^{-1}\,.$ (20)
@@
This corresponds to the Born approximation, which involves using the incident field as the driving field at each point in the scattering device, rather than the total EM field. This approximation is valid when the scattered field is small relative to the incident field on the scattering device. Notably, this result is equivalent to removing the feedback loop in Fig. 4. Depending on the specific technology being analyzed, the use of the approximation (20) instead of (19) should be considered with caution, as it may fail to capture important physical phenomena. An example of this is the Floquet modes in periodic metamaterials, which correspond to spurious reflections in undesired directions when used as a RIS [32, 15].
@@
In conclusion, the ESP problem can be framed as finding the optimal configuration parameter set $\bm{\theta}$ that best approximates a given objective end-to-end processing functionality $\mathbf{H}_{\text{o}}$} according to a specific criterion. This can be formulated as the following minimization problem:
$\hat{\bm{\theta}}=\arg\min_{\bm{\theta}}\|\mathbf{G}_{\text{R}}\,{\mathbf{C}}(\bm{\theta})\,\mathbf{G}_{\text{T}}-{\mathbf{H}_{\text{o}}}\|_{\mathrm{F}}\,$, (21)
@@
subject to a constraint on the radiated power, where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm. Depending on the adopted technology, additional constraints should be incorporated into the minimization problem to account for the specific structure and symmetries of the matrix ${\mathbf{D}}(\bm{\theta})$ and hence ${\mathbf{C}}(\bm{\theta})$. This underscores the need for tailored optimization algorithms based on physically consistent models. In this regard, specific examples will be provided in Section V-B.
@@ IV Fundamental Limits
As described in Section III-B, the number of DoF of the receiver space $\mathcal{S}_{\text{R}}$ represents the minimum number of parameters sufficient to approximate any element ${\mathsf{E}}({\mathbf{r}})$ of the signal space $\mathcal{Y}={\mathbb{L}}^{2}(\mathcal{S}_{\text{R}})$, defined in $\mathcal{S}_{\text{R}}$, up to arbitrary precision. Therefore, when ${\mathsf{E}}({\mathbf{r}})$ is the result of wave processing by the reconfigurable scattering device, a fundamental question is to determine the signal processing capability of the device in approximating the desired EM response, i.e., the number of DoF the device can manage. The ultimate limit in terms of the number of DoF of the scattering device can be obtained supposing that, through a proper configuration of the device and illumination from the source space, it is possible to generate any element ${\mathsf{J}}({\mathbf{r}})\in\mathcal{X}$, where $\mathcal{X}={\mathbb{L}}^{2}({\cal{S}})$ is the signal space of the induced current distributions on $\cal{S}$}. In the following, we show that even in the case of an unbounded space (infinite receiver’s size), the number of DoF is limited by the size of the scattering device compared to the wavelength. This number of DoF is further limited when finite source and receiver spaces are considered. In any case, these theoretical limits have to be intended as upper bounds because the particular technology considered might pose further constraints.
@@ IV-A DoF in the Unbounded Space
In the following, we revisit the main results regarding the number of DoF for the radiating EM field in an unbounded space, given a current density ${\mathsf{J}}({\mathbf{r}})$ defined within a limited space $\cal{S}$}, using signal processing arguments.
@@ IV-A1 DoF for 1D Structures
In this case, $\cal{S}$} is a line segment of length $L$, with coordinate $r_{x}\in[-L/2,L/2]$. Without loss of generality, we consider vertical EM waves polarization. Suppose $^{\mathsf{J}}({\mathbf{r}})=J(r_{x})\,\delta(r_{y})\,\delta(r_{z})\,{\hat{\mathbf{y}}}$ defined in $\cal{S}$}. It is ${\widetilde{\mathsf{J}}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{J}}({\mathbf{r}})\right]=\mathcal{F}\left[J(r_{x})\right]{\hat{\mathbf{y}}}={\widetilde{J}}(\kappa_{x})\,{\hat{\mathbf{y}}}={\widetilde{\mathsf{J}}}(\kappa_{x})$, where ${\widetilde{J}}(\kappa_{x})=\mathcal{F}\left[J(r_{x})\right]$}. It is well-known that every function $J(r_{x})$ limited in [L/2,L/2][-L/2,L/2] can be represented as an infinite series expansion of Fourier basis functions (i.e., modes) [15]
$J(r_{x})=\sum_{n=-\infty}^{\infty}j_{n}\frac{1}{\sqrt{L}}\text{rect}\left({\frac{r_{x}}{L}}\right)\,e^{\jmath k_{n}\,r_{x}}\,$, (22)
@@
${\widetilde{J}}(\kappa_{x})=\sum_{n=-\infty}^{\infty}j_{n}\,\sqrt{L}\,\text{sinc}\left({\frac{L}{2\pi}\left(\kappa_{x}-k_{n}\right)}\right)\,$, (23)
being $\text{sinc}\left({x}\right)=\sin(\pi x)/(\pi x)$. Note that, according to (7), only the modes for which $|k_{n}|\leq k_{0}=2\pi/\lambda$} radiate. Then, only $2L/\lambda$ terms in (23) fall in the visible region. [^5] More precisely, we assume that $L\gg\lambda$; otherwise, the tails of certain sinc functions centered in the non-visible region could significantly contribute to the visible region. Being the wavenumber support limited (i.e., ${\widetilde{J}}(\kappa_{x})$ band limited in $k_{0},k_{0}]$), we can say that any EM radiated field that can be generated by a current line source of length $L$ can be fully represented by sampling it in the wavenumber domain at frequencies $k_{n}=2n\pi/L$ for $|n|\leq L/\lambda$. Thus, the required number of samples leads to the number of DoF in the unbounded space ${N_{\text{u}}}=\frac{2L}{\lambda}$ (per polarization). To draw a parallel, when beamforming is operated using $N_{\text{a}}$} half-wavelength spaced antennas arranged as a uniform linear array (ULA), ${N_{\text{u}}}={N_{\text{a}}}$ orthogonal beams (i.e., $2L/\lambda$ when the antenna spacing is $\lambda/2$) can be realized. These can be adopted to reach users in every possible angular direction (e.g., for initial access). These beams correspond to the set of orthogonal transmission directions that can be spanned by the ULA considering the classical DFT-based codebook of beamforming vectors given by [3]
$\frac{1}{\sqrt{{N_{\text{a}}}}}\begin{bmatrix}1,\,\,e^{-\frac{\jmath 2\pi n}{N}},\,\,e^{-\frac{\jmath 2\pi 2n}{N}},\,\,\ldots,\,\,e^{-\frac{\jmath 2\pi({N_{\text{a}}}-1)n}{N}}\end{bmatrix}^{{\mathrm{T}}}=\frac{1}{\sqrt{{N_{\text{a}}}}}\begin{bmatrix}1,\,\,e^{-\frac{\jmath k_{n}\lambda}{2}},\,\,e^{-\frac{\jmath 2k_{n}\lambda}{2}},\,\,\ldots,\,\,e^{-\frac{\jmath({N_{\text{a}}}-1)k_{n}\lambda}{2}}\end{bmatrix}^{{\mathrm{T}}}\,.$ (24)
The different vectors correspond to the transmission directions towards angles $\gamma_{n}=\arcsin{\frac{2n}{{N_{\text{a}}}}}$, for $n=\pm 1,\pm 2,\ldots,\left\lfloor\frac{{N_{\text{a}}}}{2}\right\rfloor$. The characteristic of these beams is that each one points towards the null directions of all the other beams. Every possible channel vector (i.e., direction) can be realized as a linear combination of the vectors in (24), which is an orthonormal basis for $\mathbb{C}^{{N_{\text{a}}}}$.